# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v2

@@ VII-A Evaluation Metrics
Performance evaluation of edge-cloud collaborative systems requires a comprehensive assessment across multiple dimensions. We categorize the relevant metrics into four primary groups: system performance, model efficiency, quality of service, and security and privacy. System performance metrics. Fig. 8 shows comprehensive evaluation spans system-level, model-level, service-level, and trust-level metrics, considering real-world constraints like network variability, compute heterogeneity, and regulatory demands.
@@ VII-A2 Model Efficiency
Efficient AI models are essential for effective deployment and operation within the resource-constrained edge-cloud continuum. Key metrics for quantifying model efficiency include computational cost, memory usage, and inference time, informing model selection and deployment strategies. Computational cost, typically measured in FLOPS [255], quantifies processing demands. Higher FLOPS values indicate greater computational requirements and potentially increased energy consumption, impacting resource-limited edge devices and overall system sustainability [256, 257]. Memory usage reflects a model’s storage footprint; large models may exceed the limited capacity of edge devices [258], necessitating compression techniques like pruning, quantization, or knowledge distillation [259]. Inference time, the duration for processing input and generating output, directly impacts application responsiveness [258]. In latency-sensitive applications like autonomous driving or real-time video analytics, minimizing inference time is crucial [17]. This optimization often involves balancing model accuracy and computational efficiency, potentially employing strategies like model partitioning across different devices [256, 260].
@@ VII-B Benchmarking Methods
Performance benchmarking of edge-cloud collaborative systems requires standardized evaluation approaches to facilitate meaningful comparisons across diverse system designs and optimization strategies. Standardized datasets and workloads, discussed in Sec. VII-B1, provide a common basis for evaluation, minimizing variability stemming from data discrepancies. Well-defined testing environments, explored in Sec. VII-B2, further enhance reliability and reproducibility by controlling hardware and software configurations, network setups, and other influential environmental factors. Sec. VII-B3 introduces performance comparison frameworks, offering structured methodologies and tools for identifying performance bottlenecks and areas for improvement in edge-cloud systems. Additionally, Sec. VII-B4 examines various evaluation methodologies, including simulation, emulation, and real-world testing, analyzing the strengths and limitations of each to guide researchers in selecting the most appropriate approach for their specific benchmarking requirements.
@@ VII-B1 Standard Datasets and Workloads
Consistent and comparable performance evaluation across different studies of ECCC systems requires the use of standard datasets and workloads [268]. Such standardization clarifies the strengths and weaknesses of various architectures, optimization techniques, and algorithms, enabling researchers to build upon existing work and accelerate the development of innovative solutions. For example, Edge AIBench [269] offers benchmarks modeling typical edge scenarios like ICU patient monitoring and autonomous vehicles, focusing on data distribution and workload collaboration across client devices, the edge, and the cloud for comprehensive end-to-end performance evaluation. Similarly, EdgeBench [270] provides a workflow-based benchmarking approach, allowing customization of workflow logic, data storage, and the distribution of workflow stages across computing tiers to tailor benchmarks to specific application scenarios. Appropriate workload selection is also crucial [17]. Specifically, workloads should reflect real-world application characteristics, including data intensity, computational complexity, and communication patterns, particularly in distributed intelligence scenarios where data and model distribution, as well as communication overhead for model training and inference, are critical factors. KheOps [271] further emphasizes reproducibility in edge-to-cloud experiments by providing a collaborative environment with an experiment repository, a notebook environment, and a multi-platform methodology, promoting transparency and rigor in performance evaluation. The availability of standard datasets and workloads also facilitates the development of new optimization techniques [272], allowing researchers to evaluate their methods on established benchmarks [260].
@@ VII-B2 Testing Environments
Robust benchmarking of edge-cloud collaborative systems mandates controlled testing environments [273]. Specifically, accurately capturing performance characteristics requires careful configuration and control due to the complex interactions among edge devices, network infrastructure, and cloud servers. This process involves meticulously configuring hardware, software, and network parameters to minimize external influences and ensure reproducibility. The diverse hardware platforms in edge deployments, from resource-constrained devices to powerful edge servers, necessitate standardized testing environments for fair comparisons across different system architectures [269]. For example, BenchFaaS and KheOps employ heterogeneous testbeds, encompassing virtual machines, Raspberry Pis, and various cloud platforms, to evaluate performance under diverse hardware and network conditions [274, 271]. Additionally, accurate measurement of key performance indicators (KPIs), such as latency, throughput, and energy consumption, requires precise control over network conditions, including bandwidth, latency, and packet loss. Standardized software configurations, including operating systems, libraries, and runtime environments, are also essential for eliminating inconsistencies and ensuring comparability across experiments. Representative benchmarks that accurately reflect real-world scenarios, encompassing data volume, processing requirements, and communication patterns, are necessary for comprehensive performance evaluation.
VII-B3 Performance Comparison Frameworks
@@
Benchmarking edge-cloud collaborative systems relies on performance comparison frameworks, which provide structured methodologies for evaluating diverse architectures and optimization techniques. Such frameworks offer a standardized approach to performance assessment, enabling researchers and practitioners to identify bottlenecks and derive best practices. Inspired by frameworks like OpenPerf for benchmarking open-source projects [275] and a cloud infrastructure evaluation framework [268], a robust framework for edge-cloud systems should encompass diverse workloads and metrics, facilitating systematic analysis of system configurations and algorithms for improved designs and optimized deployments. Specifically, a comprehensive framework should define a multifaceted set of evaluation metrics encompassing system performance (throughput, latency, scalability), model efficiency, quality of service, and security and privacy [276]. Additionally, it should specify standardized datasets and workloads representative of real-world applications [273], alongside guidelines for establishing controlled testing environments for reproducible results [277]. Finally, incorporating tools and methodologies for result analysis, including statistical analysis and visualization techniques [17], allows for identifying performance bottlenecks and understanding design trade-offs.
VII-B4 Evaluation Methodologies
Benchmarking ECCC systems relies on diverse methodologies, each with inherent strengths and limitations. These methodologies are broadly categorized into simulation, emulation, and real-world testing. Simulation creates a virtual model of the entire system, encompassing hardware, network characteristics, and workloads [17]. This approach offers cost-effectiveness and flexibility, enabling exploration of various scenarios and configurations without physical deployment. However, reliance on simplified models and assumptions may not accurately reflect real-world complexities, potentially leading to discrepancies between simulated and actual performance. Specifically, accurately modeling dynamic network conditions or heterogeneous edge devices remains challenging. Emulation deploys real software components on virtualized infrastructure, often incorporating realistic network emulation tools [126, 278]. Consequently, this approach bridges the gap between simulation and real-world testing by providing a more realistic yet controlled environment. Researchers can emulate network characteristics like latency, bandwidth, and packet loss using tools like NetEm or eBPF-based solutions [278], enabling more accurate performance assessments. Nevertheless, emulation still introduces virtualization overhead and may not fully capture hardware-software interactions; additionally, scaling network emulation for large deployments presents significant challenges [278]. Real-world testing deploys the system on physical infrastructure [268, 98], offering the most accurate performance assessment by capturing the full system complexity and real-world interactions. This approach is essential for validating simulation and emulation findings and demonstrating solution practicality. However, real-world testing can be expensive, time-consuming, and difficult to control, especially for large-scale distributed systems. Ensuring reproducibility and replicability also remains challenging due to variations in hardware, software, and environmental factors [271]. Tools and platforms like KheOps [271] aim to facilitate reproducible experiments across different infrastructures, partially addressing these challenges.
@@ VII-C System Design Evaluation
We analyze the performance of various ECCC paradigms by comparing different architectural approaches and exploring their strengths and weaknesses across diverse application scenarios. Additionally, we examine the effectiveness of optimization techniques, such as load balancing, resource allocation, and task scheduling, for enhancing system performance. Inherent trade-offs in ECCC, including the balance between performance and energy consumption, latency and accuracy, and security and usability, are also discussed. Specifically, Sec. VII-C1 compares different edge-cloud architectures and their suitability for various applications. Sec. VII-C2 then analyzes how various optimization techniques impact system performance and efficiency. Following this, Sec. VII-C3 explores the inherent trade-offs in edge-cloud systems, providing insights into design considerations.
VII-C1 Comparison of Different Architectures
Edge-cloud collaborative architectures balance centralized and decentralized processing, impacting performance across various applications. Cloud-centric architectures, relying heavily on cloud resources, offer scalability [279], but may introduce latency, especially in real-time applications [253]. In contrast, edge-centric architectures, such as Armada [279], prioritize local processing, minimizing latency but facing limitations in computational resources and data synchronization. This trade-off is further explored in [98], which highlights the performance benefits of localized edge processing and the scalability of cloud processing for synchronized applications. Hybrid architectures combine both paradigms. For example, AppealNet [120] uses predictive offloading to optimize accuracy and efficiency, while EdgeMatrix [280] employs a networked multi-agent actor-critic algorithm for resource management and service orchestration. Similarly, ACE [45] advocates for a unified platform to manage edge and cloud resources transparently. Architectural choices also influence model optimization. Model compression techniques suit resource-constrained edge devices [120]. Federated learning enables collaborative training across distributed edge devices while preserving privacy [45]. The SPEC-RG reference architecture [17] provides a framework for analyzing these techniques, and reinforcement learning approaches are increasingly used for dynamic resource management in complex edge-cloud systems [111].
VII-C2 Analysis of Optimization Techniques
Load balancing, resource allocation, and task scheduling are essential optimization techniques in edge-cloud collaborative systems for enhanced performance and efficiency [281]. Load balancing distributes workloads across edge devices and the cloud to mitigate congestion and improve resource utilization, as exemplified by the two-timescale optimization framework in [125] for balancing edge and cloud loads. Similarly, [123] introduces a collaborative approach for service placement and task scheduling, incorporating both edge-cloud and edge-edge cooperation for improved task offloading. Resource allocation strategies optimize the use of computational resources, bandwidth, and storage capacity. Dynamic resource allocation adapts to real-time demands and availability [167], while collaborative resource management coordinates allocation across multiple edge nodes and the cloud [19], with reinforcement learning emerging as a promising approach for adaptive resource management [53]. Finally, task scheduling determines task execution order and timing to minimize latency and improve system performance. Latency-aware and energy-efficient offloading mechanisms prioritize tasks with deadlines or minimize energy consumption [125], while dependency-aware offloading considers task dependencies for complex applications [123].
VII-C3 Trade-off Analysis
Optimizing ECCC systems requires careful consideration of various trade-offs inherent in their design. A primary concern is balancing performance, often measured by throughput and processing speed, against energy consumption [133]. While deploying computationally intensive tasks at the edge minimizes latency and improves responsiveness, it can strain the limited energy resources of edge devices; conversely, offloading to the cloud conserves edge energy but introduces network latency and potential bandwidth bottlenecks, particularly affecting latency-sensitive applications [282]. Achieving an optimal balance necessitates sophisticated resource allocation and task offloading strategies [134]. Another key trade-off lies between latency and accuracy, especially in AI-driven applications. Complex, accurate AI models demand substantial computational resources, potentially increasing latency, especially on edge devices, whereas simpler models offer lower latency but potentially reduced accuracy. Model compression techniques, such as pruning, quantization, and knowledge distillation, aim to mitigate this trade-off [140], but the optimal balance remains application-dependent. For example, real-time applications like autonomous driving prioritize minimal latency [253], potentially accepting slightly lower accuracy, while applications like medical diagnosis prioritize accuracy even at the cost of increased latency. Additionally, security and usability present another trade-off. Robust security measures, like encryption and authentication, protect data integrity but can introduce complexity and overhead, impacting usability. Specifically, complex authentication procedures, while enhancing security, can increase user login time, and data encryption adds computational overhead, potentially affecting application performance. The distributed nature of these systems further complicates security, particularly in managing trust and secure communication [53].
@@ VIII-A Technical Challenges
Edge-cloud system faces key technical hurdles that impede its widespread adoption and optimal performance. As illustrated in Fig. 9, the deployment of AI in ECCC systems faces multifaceted technical challenges, including model optimization under resource constraints, integration of heterogeneous hardware/software, overcoming performance bottlenecks, and addressing domain-specific concerns such as real-time processing and scalability. Model optimization for the unique resource constraints of edge-cloud environments requires careful consideration of the trade-offs between accuracy, efficiency, and scalability. Additionally, integrating diverse hardware and software components necessitates effective heterogeneity management, seamless interoperability, and standardized protocols. Performance bottlenecks, including communication overhead, computational and storage limitations, and the demands of real-time processing, further challenge the efficiency and responsiveness of these distributed systems. Addressing these obstacles is crucial for unlocking the full potential of this promising paradigm.
@@ VIII-A1 Model Optimization Issues
Optimizing AI models for deployment in edge-cloud collaborative systems presents significant challenges in balancing model accuracy with computational efficiency given the resource constraints of edge devices [283]. Specifically, complex and accurate models often demand substantial resources, unsuitable for edge deployment, while smaller models may sacrifice accuracy for performance [283]. Resource limitations across the edge-cloud continuum, including limited processing power, memory capacity, and energy availability, especially for IoT devices [133, 284], further complicate model deployment. Additionally, scaling AI models across distributed systems introduces complexities in maintaining performance and efficiency with growing system size and data volumes [12]. Such scaling challenges are exacerbated by communication overhead, resource contention, and device heterogeneity [285, 286]. Therefore, effective model optimization necessitates techniques like pruning, quantization, knowledge distillation [39], and resource-aware strategies like dynamic model selection and offloading [53, 30], while also considering energy-aware scheduling [287] and adaptive model partitioning [211] to address resource constraints and scalability limitations in these dynamic environments. The growing use of LLMs further emphasizes the need for efficient inference and deployment techniques on resource-constrained edge devices [288]. For example, Shoggoth [289] utilizes online knowledge distillation to enhance accuracy while managing data drift and offloading computationally intensive tasks.

### v4

@@ VII-A Evaluation Metrics
Comprehensive performance assessment of collaborative computing systems requires multidimensional evaluation that addresses the unique characteristics of distributed computing paradigms. Unlike traditional centralized systems, ECCC performance evaluation must consider the complex interplay between edge proximity and cloud scalability, heterogeneous resource constraints, and dynamic network conditions. We categorize the relevant metrics into four primary groups tailored for edge-cloud environments: system performance, model efficiency, quality of service, and security and privacy. As illustrated in Fig. 7, comprehensive evaluation spans system-level, model-level, service-level, and trust-level metrics, considering real-world constraints unique to edge-cloud deployments such as network variability, compute heterogeneity, device mobility, and multi-tier resource orchestration.
Performance evaluation in ECCC environments must address several unique challenges: (1) Cross-tier latency analysis, measuring end-to-end response times that include network transmission delays, edge processing times, and cloud computation overhead; (2) Resource utilization efficiency, evaluating how effectively computational, storage, and communication resources are distributed and utilized across the edge-cloud continuum; (3) Adaptation overhead, quantifying the costs associated with dynamic model partitioning, task migration, and real-time workload redistribution; (4) Collaborative intelligence effectiveness, assessing the performance gains achieved through edge-cloud cooperation compared to standalone edge or cloud-only deployments; (5) Fault tolerance and resilience, measuring system robustness against edge device failures, network partitions, and cloud service interruptions. Specialized considerations necessitate specialized evaluation frameworks that capture the distributed, heterogeneous, and dynamic nature of edge-cloud systems while providing meaningful comparisons with traditional computing paradigms.
@@ VII-A2 Model Efficiency
Efficient AI models are essential for effective deployment and operation within the resource-constrained edge-cloud continuum. Key metrics for quantifying model efficiency include computational cost, memory usage, and inference time, informing model selection and deployment strategies. Computational cost, typically measured in FLOPS [223], quantifies processing demands. Higher FLOPS values indicate greater computational requirements and potentially increased energy consumption, impacting resource-limited edge devices and overall system sustainability [224, 225]. Memory usage reflects a model’s storage footprint; large models may exceed the limited capacity of edge devices [226], necessitating compression techniques like pruning, quantization, or knowledge distillation [70]. Inference time, the duration for processing input and generating output, directly impacts application responsiveness [226]. In latency-sensitive applications like autonomous driving or real-time video analytics, minimizing inference time is crucial [20]. Efficiency optimization often involves balancing model accuracy and computational efficiency, potentially employing strategies like model partitioning across different devices [227].
@@ VII-B Benchmarking Methods
Performance benchmarking of edge-cloud collaborative systems requires standardized evaluation approaches to facilitate meaningful comparisons across diverse system designs and optimization strategies.
TABLE VI: Comparison of edge-cloud benchmarking tools and platforms.
Tool Focus Area Key Features Advantages Limitations Use Case Scenarios
Edge AIBench [236] End-to-end edge AI performance Models typical edge scenarios (ICU monitoring, autonomous vehicles), comprehensive client-edge-cloud evaluation Realistic application scenarios Comprehensive performance metrics Multi-tier evaluation coverage Limited customization options Fixed scenario constraints Domain-specific focus Healthcare monitoring, autonomous driving, real-time inference applications
EdgeBench [237] Workflow-based edge computing Customizable workflow logic, configurable data storage, flexible stage distribution across computing tiers High customization flexibility Workflow-centric approach Adaptable to diverse applications Complex configuration process Requires workflow expertise Limited standardized scenarios Custom application workflows, multi-stage processing pipelines, research environments
BenchFaaS [238] Serverless edge computing Heterogeneous testbeds (VMs, Raspberry Pi, cloud platforms), FaaS performance evaluation Serverless focus Heterogeneous platform support Function-level granularity Limited to FaaS paradigm Cold start overhead Platform-dependent results Serverless applications, function-as-a-service deployment, microservices evaluation
KheOps [239] Reproducible edge-to-cloud experiments Collaborative environment, experiment repository, notebook interface, multi-platform methodology Strong reproducibility support Collaborative features Multi-platform compatibility Learning curve for new users Resource-intensive setup Limited real-time capabilities Research collaboration, experimental reproducibility, academic studies
OpenPerf [240] Open-source project benchmarking General performance evaluation framework, standardized metrics, community-driven development Open-source transparency Community contributions Standardized approach Not edge-cloud specific Generic rather than specialized Limited edge device support General performance testing, open-source projects, baseline comparisons
Table VI provides a comprehensive comparison of existing benchmarking tools and their respective strengths and limitations across different edge-cloud deployment scenarios. Standardized datasets and workloads, discussed in Sec. VII-B1, provide a common basis for evaluation, minimizing variability stemming from data discrepancies. Well-defined testing environments, explored in Sec. VII-B2, further enhance reliability and reproducibility by controlling hardware and software configurations, network setups, and other influential environmental factors. Performance comparison frameworks and evaluation methodologies, including simulation, emulation, and real-world testing approaches, offer structured methodologies and tools for identifying performance bottlenecks and areas for improvement in edge-cloud systems, with detailed analysis provided in the Supplementary Material.
The choice of appropriate benchmarking tools depends on several factors including application domain, system architecture, and evaluation objectives. Edge AIBench [236] excels in scenarios requiring realistic application modeling with comprehensive end-to-end performance analysis, making it particularly suitable for healthcare and autonomous driving applications where domain-specific workloads are crucial. EdgeBench [237] offers superior flexibility for research environments and custom applications that require tailored workflow configurations, though it demands significant expertise in workflow design and configuration. BenchFaaS [238] provides specialized evaluation capabilities for serverless edge computing paradigms, offering detailed insights into function-level performance characteristics but with limited applicability outside the FaaS domain. KheOps [239] stands out for collaborative research environments where reproducibility and experiment sharing are paramount, though it may require substantial initial setup overhead. The selection process should consider factors such as workload representativeness, customization requirements, evaluation scope, and available expertise. Additionally, the benchmarking tool should align with the specific performance metrics of interest, whether focusing on latency, throughput, energy efficiency, or comprehensive system-level analysis.
@@ VII-B1 Standard Datasets and Workloads
Consistent and comparable performance evaluation across different studies of ECCC systems requires the use of standard datasets and workloads [241]. Unified standardization clarifies the strengths and weaknesses of various architectures, optimization techniques, and algorithms, enabling researchers to build upon existing work and accelerate the development of innovative solutions. For example, Edge AIBench [236] offers benchmarks modeling typical edge scenarios like ICU patient monitoring and autonomous vehicles, focusing on data distribution and workload collaboration across client devices, the edge, and the cloud for comprehensive end-to-end performance evaluation. Similarly, EdgeBench [237] provides a workflow-based benchmarking approach, allowing customization of workflow logic, data storage, and the distribution of workflow stages across computing tiers to tailor benchmarks to specific application scenarios. Appropriate workload selection is also crucial [20]. Specifically, workloads should reflect real-world application characteristics, including data intensity, computational complexity, and communication patterns, particularly in distributed intelligence scenarios where data and model distribution, as well as communication overhead for model training and inference, are critical factors. KheOps [239] further emphasizes reproducibility in edge-to-cloud experiments by providing a collaborative environment with an experiment repository, a notebook environment, and a multi-platform methodology, promoting transparency and rigor in performance evaluation. The availability of standard datasets and workloads also facilitates the development of new optimization techniques [242], allowing researchers to evaluate their methods on established benchmarks [227].
@@ VII-B2 Testing Environments
Robust benchmarking of edge-cloud collaborative systems mandates controlled testing environments [243]. Specifically, accurately capturing performance characteristics requires careful configuration and control due to the complex interactions among edge devices, network infrastructure, and cloud servers. Achieving reliable benchmarking results necessitates meticulous configuration of hardware, software, and network parameters to minimize external influences and ensure reproducibility. The diverse hardware platforms in edge deployments, from resource-constrained devices to powerful edge servers, necessitate standardized testing environments for fair comparisons across different system architectures [236]. For example, BenchFaaS and KheOps employ heterogeneous testbeds, encompassing virtual machines, Raspberry Pis, and various cloud platforms, to evaluate performance under diverse hardware and network conditions [238, 239]. Additionally, accurate measurement of key performance indicators (KPIs), such as latency, throughput, and energy consumption, requires precise control over network conditions, including bandwidth, latency, and packet loss. Standardized software configurations, including operating systems, libraries, and runtime environments, are also essential for eliminating inconsistencies and ensuring comparability across experiments. Representative benchmarks that accurately reflect real-world scenarios, encompassing data volume, processing requirements, and communication patterns, are necessary for comprehensive performance evaluation.
@@
Performance comparison frameworks offer structured methodologies for systematic evaluation of diverse ECCC architectures and optimization techniques, providing standardized approaches to performance assessment that enable identification of bottlenecks and derivation of best practices [240]. Comprehensive frameworks encompass multifaceted evaluation metrics including system performance (throughput, latency, scalability), model efficiency, quality of service, and security considerations [244], alongside standardized datasets and workloads representative of real-world applications [243]. Essential components include guidelines for establishing controlled testing environments for reproducible results [245] and tools for result analysis incorporating statistical and visualization techniques to identify performance bottlenecks and understand design trade-offs [20]. Evaluation methodologies span simulation for cost-effective scenario exploration, emulation for realistic yet controlled assessment using network emulation tools [128], and real-world testing for comprehensive validation, each offering distinct advantages in capturing system complexities while addressing specific evaluation requirements. Detailed performance comparison frameworks, systematic evaluation approaches, and comprehensive methodology analysis are provided in LABEL:sec:supp-frameworks and LABEL:sec:supp-methodologies of the Supplementary Material.
@@ VII-C System Design Evaluation
Comprehensive performance assessment across ECCC paradigms requires multi-dimensional analysis through architectural comparisons, optimization evaluations, and systematic trade-off analysis [246, 123, 247]. Architectural configurations reveal how cloud-centric solutions prioritize scalability versus edge-centric approaches like Armada emphasizing minimal latency, while hybrid systems including AppealNet and EdgeMatrix leverage predictive offloading and multi-agent algorithms for optimal balance. Additionally, optimization technique analysis examines load balancing effectiveness through two-timescale frameworks, collaborative edge-edge cooperation, and reinforcement learning-based adaptive management demonstrating superior performance in distributed scenarios [248, 249, 126]. Furthermore, trade-off analysis addresses design considerations balancing performance versus energy consumption, latency-accuracy optimization, and security-usability considerations [142, 127, 62]. Detailed architectural comparisons, optimization analysis, and trade-off methodologies are provided in LABEL:sec:supp-system-design-eval of the Supplementary Material.
@@ VIII-A Technical Challenges
As illustrated in Fig. 8, AI-driven ECCC systems face multifaceted technical challenges spanning model optimization under resource constraints, integration of heterogeneous hardware/software, performance bottlenecks, and addressing domain-specific concerns such as real-time processing and scalability. Addressing the interconnected nature of challenges requires comprehensive solutions that balance accuracy, efficiency, and scalability while ensuring seamless interoperability and overcoming practical deployment barriers.
@@ VIII-A1 Model Optimization Issues
Model optimization for edge-cloud deployment fundamentally challenges the accuracy-efficiency trade-off, where complex models demand substantial resources unsuitable for edge deployment while smaller models sacrifice accuracy [250]. Resource limitations including processing power, memory capacity, and energy availability, particularly for IoT devices [142, 251], compound deployment complexities alongside scaling challenges from communication overhead, resource contention, and device heterogeneity [252, 253, 15]. Effective optimization necessitates integrated approaches combining pruning, quantization, knowledge distillation [48] with resource-aware strategies like dynamic model selection, offloading [62, 39], energy-aware scheduling [254], and adaptive model partitioning [195]. LLMs further intensify optimization challenges, requiring specialized inference techniques for resource-constrained environments, with solutions like Shoggoth [255] demonstrating online knowledge distillation to enhance accuracy while managing data drift and computational offloading [256].